Grade 12Trigonometry

The angle of elevation of top of vertical tower from a point P on the ground is A. After 2metres from P towards foot the elevation change to D. Then the height of tower is

Srujith
7 Years agoGrade 12

1 Answer

Geo Chacko Benny
7 Years ago
let a be then distance between the foot of the tower and the point(say P) on then ground from which angle of elavation is A and
height of tower = h
then tanA=h/a.....(i)
now the distance between the point(say Q) on the ground from which angle of elavation is D
then tanD=h/(a-2) ….(ii)            [since he walks 2 metres towards the tower ….. base length becomes a-2]
now eqn(i)*eqn(ii)
tanAtanD=h^2/[a(a-2)]
from this we get h^2=tanAtanD[a(a-2)]
taking root on both side we get the height of the tower 
i am a noobie the answer may be wrng